# SpikeReveal: Unlocking Temporal Sequences from Real Blurry Inputs with Spike Streams

### v4

@@ 1 Introduction
Traditional cameras, constrained by their exposure-based imaging mechanism, often produce blurry images when capturing fast-moving objects or during camera movement throughout the exposure process [20, 33]. While these blurry images lose significant details, the ability to recover dynamic motion trajectories from the static blurry input becomes critically important. However, the inherent challenge lies in the limited motion features available within blurry frames, leading to potential ambiguities such as multiple motion trajectories corresponding to the same blurry input. This is exemplified by scenarios where two objects move along the same trajectory but in opposite directions [24, 25, 36], rendering the task of motion deblurring ill-posed. Recent advancements in learning-based approaches [39, 13, 49] seek to address this challenge by establishing direct mappings from blurry inputs to sharp sequences in a supervised learning manner. Despite these efforts, traditional cameras struggle to capture fine details in high-speed motion due to their exposure constraints, thus limiting the effectiveness of these methods in scenarios not covered by the training datasets.
@@
In recent years, neuromorphic cameras [9, 11], leveraging their ultra-high temporal resolution and high dynamic range, have found widespread use in many fields, including computer vision and robotics. These cameras, including event and spike cameras, are distinguished by their ability to produce high temporal resolution outputs directly tied to changes in light intensity. Specifically, event cameras generate events in areas where light intensity changes [4], while spike cameras capture the absolute brightness of the scene at each pixel, offering a stream of spikes as output [11]. This distinctive feature endows spike cameras with a significant advantage [50, 48, 51, 5, 40, 6] in capturing and recovering sharp texture from scenes with rapid motion.
@@
Recent studies [2, 8] have explored the potential of RGB-Spike fusion, i.e., harnessing the strengths of both traditional and spike cameras to reconstruct sharp sequences from blurry inputs. However, their frameworks are constrained within the supervised learning paradigm, which necessitates extensive datasets comprising pairs of blurry and sharp images, as well as spike sequences. While synthetically acquiring such paired data, as demonstrated in previous studies [8, 4, 25], is feasible, collecting them in real-world scenarios presents the following challenges: (1) high-speed cameras are prohibitively expensive and not readily deployable in many settings; (2) spatial-temporal calibration between spike cameras and high-speed RGB cameras complicates the data collection process. These problems render the fine-tuning of supervised methods on real-world datasets challenging, further leading to their performance deterioration in such environments as shown in Fig. 1. The resulting degradation in image quality, manifesting as color distortion, brightness inconsistency, and inaccurate texture restoration, is mainly caused by the disparity between synthetic and real-world datasets, especially in terms of the density of spike stream, spike generation mechanism, and blurry image generation. Moreover, the effectiveness of supervised methods is inherently limited by the ground truth sequences created through motion analysis interpolation algorithms [23, 12], which inherently differs from the real-world scene and thus affects the model’s generalization ability.
@@ 2 Related Work
Spike Camera. The spike camera [42, 47, 46, 44, 38], inspired by the primate retina, stands apart from conventional cameras with its ability to generate synchronous spike streams for each pixel at extremely low latency. This distinct feature provides significant advantages in various applications such as high-speed imaging [50, 48, 51, 5, 40, 6], optical flow estimation [43], object detection [45], depth estimation [32], motion deblurring [8], and occlusion removal [31].
Spike-guided Motion Deblurring. While the spike camera boasts an ultra-high temporal resolution, its development is currently impeded by the low spatial resolution. Additionally, the single-channel output from the spike camera restricts previous methods from recovering the image color information. To address these issues, a promising approach is establishing an RGB-Spike hybrid imaging system [2]. The binocular system achieves the multi-modality fusion of High-spatial/Low-temporal RGB blurry input and High-temporal/Low-spatial spike stream, thereby also serving as a spike-guided motion deblurring method [8]. However, to the best of our knowledge, existing spike-guided deblurring methods [2, 8] predominantly rely on supervised training on synthetic datasets. This reliance results in significant performance degradation when these methods are evaluated in real-world scenarios due to the domain discrepancies between synthetic and real datasets as illustrated in Fig. 1.
Event-based Motion Deblurring. Event camera [22] can asynchronously generate events that record log-intensity changes at the pixel level with minimal latency, which contains a rich set of motion features beneficial for motion deblurring tasks. Numerous supervised methods [17, 4, 3, 17, 24, 25] have been proposed to learn the mapping from the blurry input, events to the sharp outcome. Despite these advancements, a major hurdle remains in obtaining real blurry-sharp image pairs for training. To overcome the domain gap between synthetic and real-world datasets, recent methods [30, 36, 37] explored the mutual constraint between the blurry image and event stream, enabling the training of networks on real-world blur datasets.
@@ 3.1 Preliminaries
Spike Camera Mechanism. Consider $\mathbf{L}(t)$ to represent the latent sharp frame at time $t$. Each pixel $p$ in the spike camera [11, 10] has an integrator that accumulates the incoming photons at a high frequency. Once the cumulative intensity exceeds a predefined threshold $C$ at time $t_{e}$, pixel $p$ emits a spike, and the accumulation of photons is reset to zero. This process can be mathematically described as follows:
@@
Problem Formulation. In traditional photography, motion blur occurs when there is relative movement between the camera and the scene during the exposure period. According to the motion blur physical model [9], the blurry image $\mathbf{B}$ can be represented as the average of the latent frame $\mathbf{L}(t)$ over the exposure $\mathcal{T}$, i.e.:
@@ 3.2 Theoretical Analysis
To this end, we have conducted a comprehensive theoretical analysis of the spike-guided motion deblurring task, which is neglected in prior learning-based motion deblur methodologies [2, 8]. For further discussion readability, we refer to Eq. 8 as the Spike-guided Deblurring Model (SDM), which is analogous to the baseline motion deblur model EDI [19] in event camera.
@@ 3.3.1 Processing Pipeline
Denoising Network. We leverage the Blind Spot Network (BSN) [14, 1, 28, 5, 15, 7] to predict the clean spike accumulation $N_{\mathcal{T}^{\prime}}$ from the input short-exposure spike stream $\mathcal{S}_{\mathcal{T}^{\prime}}$. The core idea of BSN is to design the blind-spot strategy that compels the convolutional layer to estimate the clean value of each pixel solely based on its surrounding pixels.
Under the premise that the spike stochastic thermal noise is independent identically distributed [41], the BSN is trained to deduce sharp spike frames from the input, with the loss function formulated as:
@@
We leverage the well-explored Enhanced Deep Super-Resolution network (EDSR) [16] as the backbone of our SR network, with the loss function formulated as follows:
@@ 3.3.2 Knowledge Distillation Framework
Lightweight Deblur Network. LDN adheres to a similar input and output pattern as previous research [8], i.e., taking the blurry input $\mathbf{B}$ and the short spike stream $\mathcal{S}_{\mathcal{T}^{\prime}}$ centered around moment $t$ as inputs, with the output being the reconstructed sharp image $\mathbf{L}_{\text{ldn}}(t)$, mathematically formulated as follows:
@@
To avoid the scenario where the LDN exactly replicates the mapping of the teacher model, we design the teacher loss $\mathcal{L}_{\text{tea}}$ based on the LPIPS [34] loss and further introduce the blur reconstruction loss. The reblur loss $\mathcal{L}_{\text{reblur}}$ measures the difference between the blurry input $\mathbf{B}$ and the re-synthesized blurry image ${\widetilde{\mathbf{B}}}$, satisfying:
@@ 4.1 Dataset
Synthetic Data. For quantitative analysis of our spike-guided motion deblurring task, we construct the synthetic dataset based on the widely employed GOPRO [18] dataset. We initially utilize the interpolation algorithm XVFI [23] to augment the video frame by interpolating additional $7$ frames between each pair of consecutive sharp images. To generate the spike stream that mimics reality closely, we downsample the interpolated video to the resolution of $320\times 180$ and simulate the spike stream based on the spike simulator [41]. To replicate real-world motion blur, we synthesize each blurry input by averaging $97$ frames from interpolated video sequences.
Real-world Data. We construct an RGB-Spike binocular system and propose the first Real-world Spike-guided Blur dataset (RSB) in this community. This system consists of a fixed-exposure RGB camera (Basler acA1920-150uc) and a spike camera [11], enabling us to capture the blurry image and the corresponding spike stream simultaneously. Further details about our RSB and the spatial-temporal calibration for our binocular system are provided in the supplementary materials.
@@
LEVS[13] $\times$ 21.155 0.601 21.155 0.601 21.155 0.601 4.97M
Motion-ETR[39] $\times$ 21.955 0.610 21.955 0.610 21.955 0.610 6.55M
BiT[49] $\times$ 23.644 0.698 23.644 0.698 23.644 0.698 11.3M
TRMD[4]+DASR[27] ✓ 27.323 0.784 21.198 0.601 18.567 0.523 19.3M
RED[30]+DASR[27] ✓ 24.456 0.741 23.178 0.674 21.942 0.608 9.76M
REFID[26]+DASR[27] ✓ 28.124 0.819 15.288 0.339 13.623 0.274 15.9M
SpkDeblurNet[8]28.307 0.834 14.406 0.299 11.621 0.202 13.4M
@@ 4.2 Experimental Results
We conduct both quantitative and qualitative comparisons of our S-SDM against state-of-the-art (SOTA) motion deblurring methods, including frame-based Motion-ETR [39], LEVS [13], video-based BiT [49], event-based TRMD [4], REFID [26], RED [30] and spike-based SpkDeblurNet [8] on the GOPRO and RSB datasets. For event-based methods, we replace the event stream with the spike stream and adopt the same input representation in these methods [30, 26, 4]. We further cascade the image super-resolution technique DASR [27] as in [37] for the deblurred sequence to overcome the modality inconsistency which is not considered in these methods. We reconstruct $7$ images from one blurry input for sequence restoration evaluation [37] as listed in Tab. 1.
Results on GOPRO. To simulate the spike density domain gap as depicted in Fig. 1, we train all supervised spike-based deblurring methods on the GOPRO dataset under spike threshold [41] $V_{th}=1$ and evaluate them on datasets with spike thresholds $V_{th}=1,2,4$. Quantitative comparison results are listed in Tab. 1. Given that the principal contribution of our S-SDM is self-supervised learning, it is foreseeable that our method might be slightly inferior to the supervised methods SpkdDeblurNet on the dataset with $V_{th}=1$. While these supervised methods deteriorate on datasets with $V_{th}\neq 1$, our method achieves great generalization benefiting from the self-supervision design. Besides, our method achieves better restoration performance than the self-supervised method RED due to our consideration of the spatial-resolution mismatch between two modalities and the designed teacher loss, which imposes a stronger constraint than the optical loss in RED.
Results on RSB. We further present visualizations of single frame and sequence reconstruction comparisons on the real-world RSB dataset as depicted in Fig. 3 and 4. Both frame-based and video-based approaches fail to replicate fine textures and detailed elements present in the blurry input. While SpkDeblurNet is capable of recovering structural details and the motion trajectory, it is deteriorated by significant noise and color distortion under conditions with lower spike density than those simulated in the GOPRO dataset. Similarly, in scenarios of higher spike stream densities, the restoration of SpkDeblurNet tends to exhibit over-exposure, resulting in brightness inconsistency compared to the blurry input. Our method addresses these challenges by finetuning on the RSB dataset, ensuring that the restored sequence aligns with the real-captured blurry input and spike stream. More comparative experiments and analyses are accessible in the supplementary material.
@@ 4.3 Ablation Study
We perform ablation experiments on the GOPRO and RSB datasets to evaluate the performance of each module within S-SDM, as well as the overall effectiveness of our distillation learning framework. In this subsection, we evaluate performance using the single middle frame for simplicity.

### v5

@@ Abstract
{NoHyper}* Equal contributors.{NoHyper}🖂 Corresponding authors.
@@ 1 Introduction
Traditional cameras, constrained by their exposure-based imaging mechanism, often produce blurry images when capturing fast-moving objects or during camera movement throughout the exposure process exposure ; Survey . While these blurry images lose significant details, the ability to recover dynamic motion trajectories from the static blurry input becomes critically important. However, the inherent challenge lies in the limited motion features available within blurry frames, leading to potential ambiguities such as multiple motion trajectories corresponding to the same blurry input. This is exemplified by scenarios where two objects move along the same trajectory but in opposite directions E-CIR ; EFNet ; evdi , rendering the task of motion deblurring ill-posed. Recent advancements in learning-based approaches motion_etr ; levs ; bit seek to address this challenge by establishing direct mappings from blurry inputs to sharp sequences in a supervised learning manner. Despite these efforts, traditional cameras struggle to capture fine details in high-speed motion due to their exposure constraints, thus limiting the effectiveness of these methods in scenarios not covered by the training datasets.
@@
In recent years, neuromorphic cameras event_review ; spikecamera , leveraging their ultra-high temporal resolution and high dynamic range, have found widespread use in many fields, including computer vision and robotics. These cameras, including event and spike cameras, are distinguished by their ability to produce high temporal resolution outputs directly tied to changes in light intensity. Specifically, event cameras generate events in areas where light intensity changes trmd , while spike cameras capture the absolute brightness of the scene at each pixel, offering a stream of spikes as output spikecamera . This distinctive feature endows spike cameras with a significant advantage tfp_tfi ; STDP_zheng ; zhu2020retina ; bsn_chen ; spk2img ; red_chen in capturing and recovering sharp texture from scenes with rapid motion.
@@
Recent studies 1000FPS ; spike_deblur have explored the potential of RGB-Spike fusion, i.e., harnessing the strengths of both traditional and spike cameras to reconstruct sharp sequences from blurry inputs. However, their frameworks are constrained within the supervised learning paradigm, which necessitates extensive datasets comprising pairs of blurry and sharp images, as well as spike sequences. While synthetically acquiring such paired data, as demonstrated in previous studies spike_deblur ; trmd ; EFNet , is feasible, collecting them in real-world scenarios presents the following challenges: (1) high-speed cameras are prohibitively expensive and not readily deployable in many settings; (2) spatial-temporal calibration between spike cameras and high-speed RGB cameras complicates the data collection process. These problems render the fine-tuning of supervised methods on real-world datasets challenging, further leading to their performance deterioration in such environments as shown in Fig. 1. The resulting degradation in image quality, manifesting as color distortion, brightness inconsistency, and inaccurate texture restoration, is mainly caused by the disparity between synthetic and real-world datasets, especially in terms of the density of spike stream, spike generation mechanism, and blurry image generation. Moreover, the effectiveness of supervised methods is inherently limited by the ground truth sequences created through motion analysis interpolation algorithms sim2021xvfi ; RIFE , which inherently differs from the real-world scene and thus affects the model’s generalization ability.
@@ 2 Related Work
Spike Camera. The spike camera, inspired by the primate retina, stands apart from conventional cameras with its ability to generate synchronous spike streams for each pixel at extremely low latency. This distinct feature provides significant advantages in various applications such as high-speed imaging tfp_tfi ; STDP_zheng ; zhu2020retina ; bsn_chen ; spk2img ; red_chen ; wgse , optical flow estimation spike_flow , object detection motion_estimation_zheng , 3D reconstruction zhang2024spikegs , depth estimation spikedepth , motion deblurring spike_deblur ; zhang2024deblur , and occlusion removal spike_sai .
Spike-guided Motion Deblurring. While the spike camera boasts an ultra-high temporal resolution, its development is currently impeded by the low spatial resolution. Additionally, the single-channel output from the spike camera restricts previous methods from recovering the image color information. To address these issues, a promising approach is establishing an RGB-Spike hybrid imaging system 1000FPS . The binocular system achieves the multi-modality fusion of High-spatial/Low-temporal RGB blurry input and High-temporal/Low-spatial spike stream, thereby also serving as a spike-guided motion deblurring method spike_deblur . However, to the best of our knowledge, existing spike-guided deblurring methods 1000FPS ; spike_deblur predominantly rely on supervised training on synthetic datasets. This reliance results in significant performance degradation when these methods are evaluated in real-world scenarios due to the domain discrepancies between synthetic and real datasets as illustrated in Fig. 1.
Event-based Motion Deblurring. Event camera eventcamera1 can asynchronously generate events that record log-intensity changes at the pixel level with minimal latency, which contains a rich set of motion features beneficial for motion deblurring tasks. Numerous supervised methods LEDIVDI ; trmd ; RES ; LEDIVDI ; E-CIR ; EFNet have been proposed to learn the mapping from the blurry input, events to the sharp outcome. Despite these advancements, a major hurdle remains in obtaining real blurry-sharp image pairs for training. To overcome the domain gap between synthetic and real-world datasets, recent methods red ; evdi ; gem explored the mutual constraint between the blurry image and event stream, enabling the training of networks on real-world blur datasets.
@@ 3.1 Preliminaries
Spike Camera Mechanism. Consider $\mathbf{L}(t)$ to represent the latent sharp frame at time $t$. Each pixel $p$ in the spike camera spikecamera has an integrator that accumulates the incoming photons at a high frequency. Once the cumulative intensity exceeds a predefined threshold $C$ at time $t_{e}$, pixel $p$ emits a spike, and the accumulation of photons is reset to zero. This process can be mathematically described as follows:
@@
Problem Formulation. In traditional photography, motion blur occurs when there is relative movement between the camera and the scene during the exposure period. According to the motion blur physical model event_review , the blurry image $\mathbf{B}$ can be represented as the average of the latent frame $\mathbf{L}(t)$ over the exposure $\mathcal{T}$, i.e.:
@@ 3.2 Theoretical Analysis
To this end, we have conducted a comprehensive theoretical analysis of the spike-guided motion deblurring task, which is neglected in prior learning-based motion deblur methodologies 1000FPS ; spike_deblur . For further discussion readability, we refer to Eq. 8 as the Spike-guided Deblurring Model (SDM), which is analogous to the baseline motion deblur model EDI EDI in event camera.
@@ 3.3.1 Processing Pipeline
Denoising Network. We leverage the Blind Spot Network (BSN) BSN ; bsn_add_1 ; bsn_add_2 ; bsn_chen ; bsn_add_3 ; atbsn to predict the clean spike accumulation $N_{\mathcal{T}^{\prime}}$ from the input short-exposure spike stream $\mathcal{S}_{\mathcal{T}^{\prime}}$. The core idea of BSN is to design the blind-spot strategy that compels the convolutional layer to estimate the clean value of each pixel solely based on its surrounding pixels.
Under the premise that the spike stochastic thermal noise is independent identically distributed zhao2022spikingsim , the BSN is trained to deduce sharp spike frames from the input, with the loss function formulated as:
@@
We leverage the well-explored Enhanced Deep Super-Resolution network (EDSR) edsr as the backbone of our SR network, with the loss function formulated as follows:
@@ 3.3.2 Knowledge Distillation Framework
Lightweight Deblur Network. LDN adheres to a similar input and output pattern as previous research spike_deblur , i.e., taking the blurry input $\mathbf{B}$ and the short spike stream $\mathcal{S}_{\mathcal{T}^{\prime}}$ centered around moment $t$ as inputs, with the output being the reconstructed sharp image $\mathbf{L}_{\text{ldn}}(t)$, mathematically formulated as follows:
@@
To avoid the scenario where the LDN exactly replicates the mapping of the teacher model, we design the teacher loss $\mathcal{L}_{\text{tea}}$ based on the LPIPS lpips loss and further introduce the blur reconstruction loss. The reblur loss $\mathcal{L}_{\text{reblur}}$ measures the difference between the blurry input $\mathbf{B}$ and the re-synthesized blurry image ${\widetilde{\mathbf{B}}}$, satisfying:
@@ 4.1 Dataset
Synthetic Data. For quantitative analysis of our spike-guided motion deblurring task, we construct the synthetic dataset based on the widely employed GOPRO GOPRO dataset. We initially utilize the interpolation algorithm XVFI sim2021xvfi to augment the video frame by interpolating additional $7$ frames between each pair of consecutive sharp images. To generate the spike stream that mimics reality closely, we downsample the interpolated video to the resolution of $320\times 180$ and simulate the spike stream based on the spike simulator zhao2022spikingsim . To replicate real-world motion blur, we synthesize each blurry input by averaging $97$ frames from interpolated video sequences.
Real-world Data. We construct an RGB-Spike binocular system and propose the first Real-world Spike-guided Blur dataset (RSB) in this community. This system consists of a fixed-exposure RGB camera (Basler acA1920-150uc) and a spike camera spikecamera , enabling us to capture the blurry image and the corresponding spike stream simultaneously. Further details about our RSB and the spatial-temporal calibration for our binocular system are provided in the supplementary materials.
@@
LEVSlevs $\times$ 21.155 0.601 21.155 0.601 21.155 0.601 4.97M
Motion-ETRmotion_etr $\times$ 21.955 0.610 21.955 0.610 21.955 0.610 6.55M
BiTbit $\times$ 23.644 0.698 23.644 0.698 23.644 0.698 11.3M
TRMDtrmd +DASRDASR ✓ 27.323 0.784 21.198 0.601 18.567 0.523 19.3M
REDred +DASRDASR ✓ 24.456 0.741 23.178 0.674 21.942 0.608 9.76M
REFIDrefid +DASRDASR ✓ 28.124 0.819 15.288 0.339 13.623 0.274 15.9M
SpkDeblurNetspike_deblur 28.307 0.834 14.406 0.299 11.621 0.202 13.4M
@@
Figure 5: Visual comparison of our S-SDM against other methods on the GOPRO dataset.
@@ 4.2 Experimental Results
We conduct both quantitative and qualitative comparisons of our S-SDM against state-of-the-art (SOTA) motion deblurring methods, including frame-based Motion-ETR motion_etr , LEVS levs , video-based BiT bit , event-based TRMD trmd , REFID refid , RED red and spike-based SpkDeblurNet spike_deblur on the GOPRO and RSB datasets. For event-based methods, we replace the event stream with the spike stream and adopt the same input representation in these methods red ; refid ; trmd . We further cascade the image super-resolution technique DASR DASR as in gem for the deblurred sequence to overcome the modality resolution inconsistency which is not considered in these methods. We reconstruct $7$ images from one blurry input for sequence restoration evaluation gem as listed in Tab. 1.
Results on GOPRO. To simulate the spike density domain gap as depicted in Fig. 1, we train all supervised spike-based deblurring methods on the GOPRO dataset under spike threshold zhao2022spikingsim $V_{th}=1$ and evaluate them on datasets with spike thresholds $V_{th}=1,2,4$. Quantitative comparison results are listed in Tab. 1 and the visual comparison is demonstrated in Fig. 5. Given that the principal contribution of our S-SDM is self-supervised learning, it is foreseeable that our method might be slightly inferior to the supervised methods SpkdDeblurNet on the dataset with $V_{th}=1$. While these supervised methods deteriorate on datasets with $V_{th}\neq 1$, our method achieves great generalization benefiting from the self-supervision design. Specifically, SpkDeblurNet tends to produce darker and blurrier reconstructions on images with high thresholds as shown in Fig. 5. Besides, our method achieves better restoration performance than the self-supervised method RED due to our consideration of the spatial-resolution mismatch between two modalities and the designed teacher loss, which imposes a stronger constraint than the optical loss in RED.
Results on RSB. We further present visualizations of single frame and sequence reconstruction comparisons on the real-world RSB dataset as depicted in Fig. 3 and 4. Both frame-based and video-based approaches fail to replicate fine textures and detailed elements present in the blurry input. While SpkDeblurNet is capable of recovering structural details and the motion trajectory, it is deteriorated by significant noise and color distortion under conditions with lower spike density than those simulated in the GOPRO dataset. Similarly, in scenarios of higher spike stream densities, the restoration of SpkDeblurNet tends to exhibit over-exposure, resulting in brightness inconsistency compared to the blurry input. This over-exposure affects the dynamic range of the reconstructed image, ultimately compromising the overall perceptual quality and uniformity of the restored sequence. Our method addresses these challenges by finetuning on the RSB dataset, ensuring that the restored sequence aligns with the real-captured blurry input and spike stream. More comparative experiments and analyses are accessible in the supplementary material.
@@ 4.3 Ablation Study
We perform ablation experiments on the GOPRO and RSB datasets to evaluate the performance of each module within S-SDM, the validity of our designed network architecture, as well as the overall effectiveness of our distillation learning framework. In this subsection, we evaluate the performance based on the single middle frame for simplicity.
@@
Table 2: Performance comparison between the SAN in GEM gem and our designed LDN.
Methods PSNR $\uparrow$ SSIM $\uparrow$ Params (M) $\downarrow$ Flops (G) $\downarrow$
SAN gem 27.283 0.773 2.36 107.84
LDN (Ours) 27.928 0.786 0.234 33.60
Network Architecture. While our designed LDN mirrors the Scale-aware Network (SAN) proposed in the GEM gem , we replace the SAN with the LDN to compare the performance difference between the two architectures as depicted in Tab. 2. In the self-supervised learning framework discussed in this paper, enhancing restoration performance primarily hinges on improving the quality of pseudo-labels rather than the network architecture itself. Despite its simple design, which consists only of convolutional layers, ResBlocks, and basic modules such as CBAM CBAM , LDN outperforms SAN in both PSNR and SSIM while requiring fewer parameters and less computation, demonstrating that LDN is both sufficient and efficient for the self-supervised spike-guided motion deblurring task.